machine A,B and C all produce the same two parts X and Y.of all the parts produced,machine A produces 60% ,machine B produced 30% and machine C produced 10% in addition,40% of the same parts made by machine A are part X,50% of the parts made by machine B are part X,70% of the parts made by machine C are part X,A part produced by this company is randomly sample and is determined to be an X part. What is the probability that the part came from machine A.
1
Expert's answer
2020-04-10T12:49:02-0400

The probability of production is given by,

"P(A) = 0.6,\\
P(B) = 0.3,\\
P(C) = 0.1."


Part "X" made by each machine is given by,

"P(X | A) = 0.4,\\
P( X | B) = 0.5,\\
P(X | C ) = 0.7."


Probability of producing part "X"

"P(X) = P(X | A) P(A) + P(X | B) P(B) + P(X | C) P(C)\\"

Substituting from above values gives,

"P(X)= 0.46"


From

"P(X | A) P(A) = P(A | X) P(X) \to

P(A | X) = P(X | A) P(A) / P(X),"


"P(A | X) = 0.521"

Answer: "P(A | X) = 0.521"